An athlete travels from point A directly to point B. If he runs 3/4 of the way, walks 1/9 of the way, and jogs the last 10 miles, what is the distance from point A to point B?
(A) 36 miles (B) 45 miles (C) 56 miles (D) 72 miles (E) none of these

Found 2 solutions by richwmiller, nerdybill:
Answer by richwmiller(17219)

3x/4+x/9+10=x
x=72
put everything over 36
27x/36+4x/36+10=x
31x/36+360/36=x
1/36(31x+360) = x

Answer by nerdybill(7384)

An athlete travels from point A directly to point B. If he runs 3/4 of the way, walks 1/9 of the way, and jogs the last 10 miles, what is the distance from point A to point B?
.
Let x = distance between A and B
then
(3/4)x + (1/9)x + 10 = x
Multiply both sides by a common denominator (say 36):
27x + 4x + 360 = 36x
31x + 360 = 36x
360 = 5x
72 miles = x